# Tax Credits and Household Behavior: The Roles of Myopic Decision-Making and Liquidity in a Simulated Economy

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 5th ACM International Conference on AI in Finance; November 14–16, 2024; Brooklyn, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
  • We consider a multi-agent economic model comprising households, a firm, central bank, and government. Our implementation expands upon a state-of-the-art financial market simulator (Byrd et al., 2019). We particularly focus on evaluating the impact of tax credits on households with different degrees of myopia and liquidity.
@@ 4.1. Multi-Agent Simulator
Our simulations are implemented using a state-of-art multi-agent simulator called ABIDES (Byrd et al., 2019), and its OpenAI Gym extension (Amrouni et al., 2021). It is a high-fidelity simulator used by practitioners and researchers in finance (Vyetrenko et al., 2020; Dwarakanath et al., 2022). Agents in ABIDES have access to their internal states, and receive information about other agents via messages. A simulation kernel handles message passing between agents and runs simulations over a specified time horizon, maintaining timestamps for all agents and the overall simulation. We adapt ABIDES to model interactions between the economic agents described in section 3, and enable MARL using its Gym extension. All agents are initialized with starting states and parameters drawn from the literature as described next (also see (Dwarakanath et al., 2024) for more details). Each agent takes actions based on its observations as in Table 1, so that agent interactions govern the evolution of the economic system to the next state. While we don’t use real data to calibrate our agent-based simulator, validation checks in (Dwarakanath et al., 2024) show it reproduces economic stylized facts. We further validate our simulator by qualitatively reproducing trends observed in real data used in the JPMorgan Chase report (Wheat et al., 2022) in section 5.2.
@@ 5.3. Proposal for tax credit distribution scheme
Previously, we saw that while household savings increased with uniform tax credits, low liquidity households still had lower savings than higher liquidity ones. This prompted us to learn a government strategy to distribute tax credits to maximize household social welfare. We consider the same set of agents as in the above experiment with the government also being a learning agent. It learns to distribute all tax revenue as credits ($\xi=1$) in a way to maximize its reward (i.e., social welfare). Learning rates are $2\times 10^{-3}$ for households, $5\times 10^{-3}$ for the firm, $5\times 10^{-3}$ for the central bank and $10^{-2}$ for the government. Discounted cumulative rewards during training are shown in Figure 6. Learned policies are tested in 1000 episodes with same regime as training, and results are plotted in the last row of Figure 5. We observe that household savings are more uniform and higher than in the previous cases, even while consumption ordering remains the same. Crucially, we see that social welfare is higher than before. This demonstrates the efficacy of the learned tax credit distribution scheme that ensures equity even when households have differing propensities to consume.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 5th ACM International Conference on AI in Finance; November 14–17, 2024; Brooklyn, NY, USAbooktitle: 5th ACM International Conference on AI in Finance (ICAIF ’24), November 14–17, 2024, Brooklyn, NY, USAdoi: 10.1145/3677052.3698599isbn: 979-8-4007-1081-0/24/11
@@ 1. Introduction
  • We consider a multi-agent economic model comprising households, a firm, central bank, and government. Our experiments utilize the economic simulator ABIDES-Economist (Dwarakanath et al., 2024), an extension of a state-of-the-art financial market simulator (Byrd et al., 2019) tailored to economic simulations. In particular, we focus on evaluating the impact of tax credits on heterogeneous households with different degrees of myopia and liquidity.
@@ 4.1. Multi-Agent Simulator
Our simulations are implemented using ABIDES-Economist, a multi-agent economic system simulator designed to facilitate reinforcement learning strategies for economic agents (Dwarakanath et al., 2024). In ABIDES-Economist, agents can access their internal states and receive information about other agents via messages. A simulation kernel handles this message passing and runs simulations over a specified time horizon, maintaining timestamps for all agents and the overall simulation. We adapt the simulator to model interactions between the economic agents described in section 3. Each agent is initialized with starting states and parameters informed by existing literature, particularly studies that examine real monetary policy (Evans and Honkapohja, 2005). These parameters are designed to reflect realistic economic behaviors, providing a robust foundation for agent decision-making. Agents act based on their observations, as outlined in Table 1, with their interactions driving the evolution of the economic system.
Although our agent-based simulator is not calibrated using real economic time series data, agent action spaces are constructed to include variations around typical values observed in real data from US Bureau of Labor Statistics (2023) e.g., labor hours chosen from $\{0,240,480,720,960\}$ where $480$ hours per quarter is the default action analogous to 40 hours per week. And, wages in \$ per hour are chosen from $\{7.25,19.65,32.06,44.46,56.87\}$ where $7.25$ is the minimum wage and $32.06$ is the default. Similarly, income tax rate choices for the government are chosen to reflect the marginal tax rates associated with the various tax brackets in the US (Internal Revenue Service, 2023). Validation checks reported in (Dwarakanath et al., 2024) demonstrate that the simulator successfully reproduces key economic stylized facts, such as the inverse relationship between firm prices and consumption, as well as the direct relationship between inflation and interest rates. We further validate our simulator by qualitatively reproducing trends observed in real data used in the JPMorgan Chase report (Wheat et al., 2022) in section 5.2.
@@ 5.3. Proposal for tax credit distribution scheme
Previously, we saw that while household savings increased with uniform tax credits, low liquidity households still had lower savings than higher liquidity ones. This prompted us to learn a government strategy to distribute tax credits to maximize household social welfare. We consider the same set of agents as in the above experiment with the government also being a learning agent. It learns to distribute all tax revenue as credits ($\xi=1$) in a way to maximize its reward (i.e., social welfare) as given in (6). Learning rates are $2\times 10^{-3}$ for households, $5\times 10^{-3}$ for the firm, $5\times 10^{-3}$ for the central bank and $10^{-2}$ for the government. Discounted cumulative rewards during training are shown in Figure 6.
Learned policies are tested in 1000 episodes with the same regime as during training, and results are plotted in the last row of Figure 5. We observe that household savings are more uniform and higher than in the previous cases, even while consumption ordering remains the same. To investigate this increase in savings of all households, we examine the price and wage set by the firm in Figure 7. Notably, when all agents learn in presence of a government policy that adapts to maximize household social welfare, the firm learns to set lower prices and higher wages than it did under a fixed government policy. Additionally, we see from the last column of Figure 5 that social welfare (as measured by the government reward) is higher than before. This demonstrates the efficacy of the learned government policy that ensures equity even when households have differing propensities to consume.
@@ 6. Conclusion
Figure 7. Distribution of average firm observables in presence of equal tax credits (first row) and presence of learned distribution of tax credits (second row).